OpenStudy (anonymous):

Solve the system. then classify the system as consistent and independent, consistent and dependent or inconsistent 2X-3Y=3 4X-5Y=9

OpenStudy (anonymous):

care to try it alone at first?

OpenStudy (anonymous):

i did i got 4,3 i just need to check it

OpenStudy (anonymous):

then put 4 in for x and 3 in for y and solve the equation to check

OpenStudy (anonymous):

i guess that's off a bit

OpenStudy (anonymous):

like this 2(4) -3(3) does that =3? and then 4(4)-5(3) does that =9? if they both don't work then try again.

OpenStudy (anonymous):

how i did what you showed me

OpenStudy (anonymous):

how i did what you showed me

OpenStudy (anonymous):

2x-3y=3 (-2)= -4x+6y=-6 4x-5y=9 ___________ y=3 2x-3*3=3 2x=12 x=6

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes its consistent independent to right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no solution inconsistent. infinite solutions con, dependent. one solution, con, indep

OpenStudy (anonymous):

yea i just did one that is no soultion